OpenStudy (anonymous):

Consider acid rain that was measured to have a pH of 2.5. How many times more acid is this acid rain than typical rain with a pH of about 6?

OpenStudy (anonymous):

pH=-log([H+]) 10^(-pH)=[H+] 10^(-2.5) --------=10^(3.5)=3162.3 times the H+ concentration 10^(-6)

OpenStudy (anonymous):

so the answer 3162.3